OpenStudy (anonymous):

Which of the following is the slope between the two points (2, 3)and (1, 4)?

OpenStudy (anonymous):

slpoe=y2-y1/x2-x1

OpenStudy (anonymous):

slope = (4-3)/(1-2)=-1

OpenStudy (anonymous):

thanks both of you